OpenStudy (anonymous):

I have a question #15 59,73,83,94,107,X What is X???

OpenStudy (anonymous):

Technically, X could be any number, depending on whatever rule you are associating to this pattern. However, an interesting solution which comes to mind is X=115. The rule used to get this solution is to take your number, say 59, and then add to it the sum of its digits. So 73 = 59+5+9, 83=73+7+3, etc., so 107+1+0+7=115. However, like I said initially, it is very important to note that X can be any value (115, pi, -6.345464, etc.) depending on whatever rule you associate to the sequence.

OpenStudy (anonymous):

thx